Session: Feb 2001 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (mauro at Sun Feb 18 19:20:15 MET 2001) Success
All scan is recorded sky clear gps- formatter= -85.3 uSec

Torun (K.Borkowski at Mon Feb 19 09:59:22 MET 2001) Minor failures
Continued problems with BBC8 (occasionaly also with BBC6). Strong winds after 15:00 (14 - 22 m/s).

Onsala (A. Polatidis at Mon Feb 19 10:55:44 MET 2001) Success
Intermittent RFI, particularly strong in the 1414-1418 MHz range. Clear weather. No other apparent problems.

Westerbork (Tony Foley at Mon Feb 19 10:58:29 MET 2001) Failure
Really bad problems with the array. This seems at present to be in the bulk delay system. This is still under repair (monday). From the start to 04:33 all telescopes were in the adder although at least 8 were not working. From 04:33 to 10:54 we used 6 telescopes,from 10:54 to 14:31 no observations were done and from 14:31 to the end we used the single dish setup (RT7)

Effelsberg (A. Kraus at Tue Feb 20 13:22:58 MET 2001) Minor failures
Several Tsys-measurements were corrupted. For few scans the fs reported that tape head one was not in position; offsets up to 3 mu occur. This is checked at the moment.

Jodrell_Bank (Alastair Gunn at Wed Feb 21 16:12:00 MET 2001) Success
No known problems (but see feedback for 18/21 cm session).

Noto (C.Trigilio at Wed Feb 28 15:33:55 MET 2001) Failure
Old version of the schedule. The message of a new version of the schedule arrived too late for problems with our e-mail server.